OpenStudy (anonymous):

At time t = π/6, the position x(t) = 10 cos t is given by the following. x(π/6) = 10 cos π/6

zepdrix (zepdrix):

Do you have a question Kori? :o

OpenStudy (anonymous):

Find the position, velocity, and acceleration of the mass at time t = π/6. In what direction is it moving at that time?

zepdrix (zepdrix):

At time t = π/6, the `position` x(t) = 10 cos t ...

zepdrix (zepdrix):

So that's our position function, yes?

zepdrix (zepdrix):

Looks like they gave us the position function evaluated at π/6 for free.\[\Large\bf\sf x(\pi/6)\quad=\quad 10\cos(\pi/6)\]

zepdrix (zepdrix):

Velocity function can be found by taking the derivative of the position function. Have you tried taking the derivative yet? :o

zepdrix (zepdrix):

\[\Large\bf\sf x(t)=10\cos t\] \[\Large\bf\sf v(t)=x'(t)\]\[\Large\bf\sf x'(t)=?\]